# Should public health policy exempt cases with low viral load from isolation during an epidemic?: a modelling study

### v1

@@ Abstract
As demonstrated during the COVID-19 pandemic, non-pharmaceutical interventions, such as case isolation, are an important element of pandemic response. The overall impact of case isolation on epidemic dynamics depends on a number of factors, including the timing of isolation relative to the onset of contagiousness for each individual instructed to isolate by public health authorities. While there is an extensive literature examining the importance of minimising the delay from exposure to direction to isolate in determining the impact of case isolation policy, less is known about how underlying epidemic dynamics may also contribute to that impact. In particular, empirical observation, and modelling studies, have shown that as an epidemic progresses, the distribution of viral loads among cases systematically changes. In principle, this may allow for more targeted and efficient isolation strategies to be implemented. Here, we developed a multi-scale agent-based model to investigate the potential for isolation strategies that account for cases viral loads to be incorporated into policy. We compare the impact and efficiency of alternative strategies in which all cases, regardless of their viral load, are required to isolate and strategies in which some cases may be exempt from isolation. Our findings show that, following the epidemic peak, the vast majority of cases identified with a low viral load are in the declining phase of their infection and so contribute little to overall contagiousness. This naturally prompts the question as to the potential public health value of discontinuing isolation for such individuals. Our numerical investigation of this ‘adaptive’ strategy shows that exempting individuals with low viral loads from isolation following the epidemic peak leads to a modest increase in new infections. Surprisingly, it also leads to a drop in efficiency, as measured by the average number of infections averted per isolated case. Our findings therefore suggest caution in adopting flexible or adaptive isolation policies. Our multi-scale modelling framework is sufficiently flexible to enable extensive numerical evaluation of more complex isolation strategies, incorporating more disease-specific biological and epidemiological features, supporting the development and evaluation of future public health pandemic response plans.
@@ 1 Introduction
One possible approach to minimising the costs associated with isolation is to make use of routinely available data on cases that may allow for more targeted application of isolation. In many jurisdictions, cases detected by qt-PCR also have an associated PCR cycle threshold (Ct) value reported to health authorities. The Ct value provides a quantitative measure of the viral load at the time of testing of each case [11, 12]. Given that an individual’s contagiousness is likely an increasing function of their viral load—an assumption for which there is a growing body of empirical evidence [13, 14]—it is therefore reasonable to consider whether those with low viral loads (high Ct values) could be exempt from a requirement to isolate, because their lower contagiousness means they are unlikely to infect other people. However, a high Ct value (low viral load) could be measured at the end of an infection (where contagiousness is decreasing) or at an early phase of an infection (where contagiousness is increasing). Failing to isolate cases with high Ct values and increasing contagiousness could reduce the effectiveness of a more targeted isolation policy. Hay et al. [15] showed that the distribution of Ct values from cases varies over the course of an epidemic wave, suggesting that the proportion of cases with high Ct values and increasing contagiousness is also likely to change during an epidemic. Therefore, the implementation of a more targeted isolation policy may have varying effectiveness at reducing infections during different phases of an epidemic.
@@
Here, we develop a multi-scale agent-based epidemic model, in which individual-level viral infection kinetics are resolved, to explore the hypothesis that exempting low viral load cases from isolation during an outbreak will have minimal effect on the final size of an outbreak but reduce the overall number of people who are required to isolate. We use the model to quantify the dynamic viral load distribution of agents throughout an unmitigated epidemic before using it to evaluate and compare several isolation strategies in which the decision to isolate is based upon viral load. We distill the problem down to its simplest form, by considering the transmission of an SIR-like pathogen in which each infected individual’s viral load is modelled using the TIV equations, An individual’s contagiousness is specified to be a sigmoidal function of their viral load. We account for test sensitivity as a function of viral load, and consider scenarios in which only one, or multiple tests, are available for individuals during their symptomatic illness. We evaluate isolation management strategies in which all individuals who test positive to infection are isolated, and alternative strategies in which those with a low viral load may be exempt from isolation. We compare isolation strategies in terms of the reduction in the final size of the epidemic compared to a baseline with no isolation, and how many individuals were isolated to achieve that reduction in overall infections.
@@ 2.1 Multi-scale model
We consider a fixed population of $N$ individuals, with no births or deaths occuring during the simulation period. Each individual in the population is represented as an agent $a_{k}$, where $k\in\{1,2,\cdots,N\}$. At each time step $t\in\{0,t_{\rm int},\dots,t_{\rm end}\}$, where $t_{\rm int}$ is the time step and $t_{\rm end}$ is the maximum simulation time, each agent is classified into one of three mutually exclusive sets: $\hat{S}(t)$ for susceptible, $\hat{I}(t)$ for infected and infectious, and $\hat{R}(t)$ for those who are removed and no longer infectious. The respective cardinalities of these sets are denoted as $S(t)$, $I(t)$, and $R(t)$, representing the total number of susceptible, infectious and recovered agents at time $t$.
@@
We define the contagiousness of an infectious agent (when $a_{k}\in\hat{I}$) as the probability of infection given contact with a susceptible individual. We model contagiousness $\tau$ units of time since the time of infection, denoted by $\beta_{k}(\tau)$, as a sigmoidal function of the agent’s viral load $V_{k}(\tau)$:
@@ 51             Calculate $S(t+t_{\rm int}),\ I(t+t_{\rm int}),\ R(t+t_{\rm int})$;
While an agent’s viral load over time is a continuous trajectory, for convenience in defining and evaluating our alternative isolation strategies, we divide the infected period of each agent into four distinct phases, defined by whether an agent’s viral load is low or high, and whether it is increasing or decreasing (Figure 1). The threshold viral load for classifying an agent as having a low or high viral load is denoted by $V^{*}$, and we choose this value such that the agent’s contagiousness $\beta=\frac{2}{3}\beta_{max}$. Combined with the sign of $dV/dt$, we compute the times of transition between the phases: $\tau_{\text{inc}}^{*}$, $\tau_{\text{peak}}$, and $\tau_{\text{dec}}^{*}$.
@@ 2.2 Testing–isolation strategies
In order for an infected individual to be isolated, they must be identified through surveillance (i.e., be recorded as a case). We consider case detection due to symptom-based testing and do not model contact tracing, testing or quarantine of contacts, nor other aspects of surveillance such as mass screening. Our model for the probability that an infected agent is identified as a case depends on the presence of symptoms, the probability $P(\rm Tested)$ per unit time of seeking a test (symptom dependent), and the sensitivity of the test (assumed to vary with viral load).
@@
Infected agents identified as cases and subject to isolation (strategy dependent) are assumed to be isolated immediately and for a fixed duration $d_{\text{iso}}$, and do not contribute to the force of infection in the population (i.e., their contagiousness is set to zero) during their isolation period. If released from isolation while still classified as an infectious agent they will once again contribute to the force of infection.
For each of the strategies, we consider alternative testing scenarios in which each agent may test only once, or in which multiple (unlimited) tests may be conducted. Under the single test strategy, each infected agent has only one opportunity for testing (regardless of test outcome). Therefore, an infectious agent with a low increasing viral load, if tested and positive (i.e., identified as a case), will not be isolated under the Isolate high strategy, even though they will soon become highly contagiousness (entering their high increasing followed by high decreasing infection phase). Under a multiple testing scenario, such an agent could test positive later in their infection and be isolated under the strategy. The same outcomes for an agent can potentially play out under the Adapativen strategy at certain times in the epidemic (depending on when the switch to only isolating cases with a high viral load is made).
@@ 2.3 Strategy evaluation
Finally, to evaluate the performance of the Adaptiven strategy in more detail, we introduce an additional metric, the relative cost, $\psi_{n}$ for the Adaptiven strategy. At best, the Adaptiven strategy should perform as well as the Isolate all strategy and at worst, perform as poorly as the Isolate high strategy. We define $\psi_{n}$ as the proportionate loss of benefit of the Adaptiven strategy between these two extremes:
@@
Accordingly, $\psi_{n}\to 0$ indicates that the effectiveness of the Adaptiven strategy is the same as for the Isolate all strategy (i.e., a cost of zero), while $\psi_{n}\to 1$ indicates that the Adaptiven strategy has lost all of the benefit of having isolated cases with a low viral load at the time of detection, with effectiveness degraded to the same value as for the Isolate high strategy.
@@ 2.4 Model parameterisations
All outputs shown in the main manuscript are generated with the “influenza like” model parameterisation. In the Supplementary Material, we present analagous outputs under the “SARS-CoV-2” model parameterisation.
For all scenarios presented (including those in the Supplementary material), we consider a population of size $N=20,000$, with $\beta_{max}=0.6$, initial conditions ${S(0)=N-10}$, ${I(0)=}10$, ${R(0)=0}$, and with all other parameters specified as per the Methods and summarised in Table 1. Our choice for $\beta_{max}$ results in an epidemiological-scale basic reproduction number of $\mathcal{R}_{0}\approx 2.12$. With these parameter choices, initial conditions, and without any intervention, it is unlikely for stochastic extinction to occur during the early phase of an epidemic.
@@ 3.1 Multi-scale dynamics
Figure 2(a) shows a single exemplar realisation of the model, which as expected displays SIR-type epidemic dynamics. For the same realisation, Figure 2(b) presents the contagiousness ($\beta$) of each infectious agent on each day, differentiated by whether the agent’s contagious is increasing (purple circles, labeled $\beta^{inc}$) or decreasing (green circles, labeled $\beta^{dec}$) on that day. Prior to the peak (around day 27), many of the infected agents are in the increasing phase of their contagiousness, whereas following the peak, most agents are in the decreasing phase of their contagiousness. Probability (PDF) and cumulative (CDF) density functions for the distribution of infected agents’ contagiousness are shown in S3 Fig..
Figures 2(c)–(f) present key aspects of our multi-scale model averaged over 20 realisations. We break the infectious population down into those with low (yellow) or high (blue) viral load (Figure 2(c)), and further, as per Figure 1, into those with a low increasing, high increasing, high decreasing and low decreasing viral load (Figure 2(e)). The contribution of these subsets of the infectious population to the overall force of infection is shown in Figures 2(d) and (f). While the proportion of infected agents with a low viral load increases throughout the outbreak, their contribution to the force of infection remains small (under 20%). Furthermore, following the epidemic peak, the vast majority of those with a low viral load are in the decreasing phase of their infection (i.e., their contagiousness is low and decreasing), suggesting that an Adaptiven strategy—in which those individuals with a low viral load at the time of detection are not be required to isolate—may prove an effective and efficient policy option.
@@ 3.2 Effectiveness and efficiency of alternative testing–isolation strategies
Figure 3 presents incidence (left panels) and cumulative incidence infection (right panels) curves for a single exemplar simulation of the agent-based model under the four strategies as defined in the Methods (No isolation, Isolate high, Isolate all and Adaptive0 with the strategy switch occurring at the peak day (i.e., $n=0$)). Results are shown under a multiple test scenario (top) and single test scenario (bottom).
@@
Unsurprisingly, the Isolate all strategy is the most effective strategy as all agents that test positive are isolated. Allowing for multiple tests results in nearly twice the number of cases averted compared to if only single tests are allowed (Figure 4(a) and (b)). The Adaptive0 strategy (with the switch enacted at the peak day) consistently yielded an intermediate effectiveness suggesting that, while the vast majority of those with a low viral load are in the decreasing phase of their infection following the peak, their collective contribution to transmission is important. A surprising result is that the efficiency—defined as the number of cases averted per person isolated—is highest for the Isolate all strategy and lowest for the Isolate high strategy. Furthermore, efficiency is higher under the single test scenario compared to the multiple test scenario (Figure 4((c) and (d)).
@@ 3.3 Sensitivity analyses: inclusion of asymptomatic infections and alternative within-host dynamics
To this point, we have assumed that all infected agents will display symptoms and that the probability of seeking a test per unit time increases when displaying symptoms. Our particular choice of parameters for those testing probabilities means that a substantial proportion of all infections are detected as cases, with corresponding strong impacts on transmission under all strategies. Here, we now allow for 30% of the infected agents to remain asymptomatic throughout their infectious period. Asymptomatic cases are assumed to have a $20\%$ chance per day of being tested, fixed throughout their infectious period. This is consistent with our assumptions for pre-symptomatic infections which have a 20% chance of testing per day. Results are shown in S4 Fig.. The trends in both effectiveness and efficiency across alternative strategies remain unchanged, with the obvious effect that with 30% of the population less likely to be tested, effectiveness is lower for any given strategy.
@@ 4 Discussion
While effective at reducing transmission, NPIs, such as case isolation, can also produce unintended health and economic costs. Therefore, it is desirable to identify pandemic response policies that would enable such measures to be targeted more selectively to reduce their social costs while maintaining their health benefits. Here, we explored the potential for isolation decisions based on routinely collected (during a pandemic) viral load data. Specifically, we explored the hypothesis that exempting low viral load cases throughout or only during the latter stages of an outbreak could reduce the overall number of people who are isolated with minimal effect on the size of the outbreak.
@@
We made a number of simplifying assumptions in our model. A quantitative exploration of alternative isolation strategies for a particular pathogen of interest would require incorporation of additional complexity. Firstly, each agent’s viral dynamics were modelled using the basic TIV model and with identical parameters (except for the initial viral load), so there is less variation across agents in our simulation than would be expected in reality. Addressing this would be technically straightforward as, at least for pathogens such as influenza and SARS-CoV-2, there are now more realistic within-host models and suitable data from which to compute population-level estimates for the joint probability distribution of within-host parameters, allowing for individual agent’s viral kinetics to be sampled. While a significant undertaking, this data-driven work would enable a more realistic population to be simulated. Secondly, our testing and isolation model is very simple. We adopted simple functional forms for both the probability of seeking a test and the probability of returning a positive result given the individual’s viral load. Empirical data on how both quantities relate to the time of exposure and symptom onset are well documented for some pathogens and could be incorporated into our multi-scale model if it were to be applied to a particular pathogen of interest. Finally, delays to isolation following testing—which depend upon aspects of the public health laboratory reporting and surveillance system—are also important, but not modelled here. Again, empirical data on those delays could be incorporated into our flexible multi-scale framework.
A natural extension to our work would be to allow for a different (shorter) duration of isolation for those with low viral loads under the Adaptiven strategy. It may be that a shorter period of isolation (compared to that for all cases identified prior to the peak of the epidemic) would prevent the vast majority of infection events, and result in an increase in efficiency, if re-defined as the infections averted per person-days of isolation (rather than per person isolated).
By developing a multi-scale model of transmission dynamics, we have been able to explore the potential impacts of alternative isolation strategies on epidemic dynamics. Our results indicate that even those with low and decreasing viral loads make a sufficient contribution to overall transmission. Any consideration of adaptive isolation policies based on viral load would require detailed evaluation using a carefully calibrated multiscale model, for which our model could provide a foundation, and likely additional surveillance streams to ensure a net positive impact of such policies.
5 Supporting information

### v2

@@ Abstract
As demonstrated during the COVID-19 pandemic, non-pharmaceutical interventions, such as case isolation, are an important element of pandemic response. The overall impact of case isolation on epidemic dynamics depends on a number of factors, including the timing of isolation relative to the onset of contagiousness for each individual instructed to isolate by public health authorities. While there is an extensive literature examining the importance of minimising the delay from exposure to direction to isolate in determining the impact of case isolation policy, less is known about how underlying epidemic dynamics may also contribute to that impact. Empirical observation and modelling studies have shown that, as an epidemic progresses, the distribution of viral loads among cases changes systematically. In principle, this may allow for more targeted and efficient isolation strategies to be implemented. Here, we describe a multi-scale agent-based model developed to investigate how isolation strategies that account for cases viral loads could be incorporated into policy. We compare the impact and efficiency of isolation strategies in which all cases, regardless of their viral load, are required to isolate to strategies in which some cases may be exempt from isolation. Our findings show that, following the epidemic peak, the vast majority of cases identified with a low viral load are in the declining phase of their infection and so contribute less to overall contagiousness. This observation prompts the question about the potential public health value of discontinuing isolation for such individuals. Our numerical investigation of this ‘adaptive’ strategy shows that exempting individuals with low viral loads from isolation following the epidemic peak leads to a modest increase in new infections. Surprisingly, it also leads to a drop in efficiency, as measured by the average number of infections averted per isolated case. Our findings therefore suggest caution in adopting such flexible or adaptive isolation policies. Our multi-scale modelling framework is sufficiently flexible to enable extensive numerical evaluation of more complex isolation strategies that incorporate more disease-specific biological and epidemiological features, supporting the development and evaluation of future public health pandemic response plans.
@@ 1 Introduction
One possible approach to minimising the social and economic costs associated with isolation is to make use of routinely available data on cases that may allow for more targeted application of isolation. In many jurisdictions, cases detected by qt-PCR also have an associated PCR cycle threshold (Ct) value reported to health authorities. The Ct value provides a quantitative measure of the viral load at the time of testing of each case [11, 12]. Given that an individual’s contagiousness is likely an increasing function of their viral load—an assumption for which there is a growing body of empirical evidence [13, 14]—it is therefore reasonable to consider whether those with low viral loads (high Ct values) could be exempt from a requirement to isolate, because their lower contagiousness means they are unlikely to infect other people. However, a high Ct value (low viral load) could be measured at the end of an infection (where contagiousness is decreasing) or at an early phase of an infection (where contagiousness is increasing). Failing to isolate cases with high Ct values and increasing contagiousness could reduce the effectiveness of a more targeted isolation policy. Hay et al. [15] showed that the distribution of Ct values from cases varies over the course of an epidemic wave, suggesting that the proportion of cases with high Ct values and increasing contagiousness is also likely to change during an epidemic. Therefore, the implementation of a more targeted isolation policy may have varying effectiveness at reducing infections during different phases of an epidemic.
@@
Here, we describe a multi-scale agent-based epidemic model in which individual-level viral infection kinetics are resolved, and use it to explore the hypothesis that exempting low viral load cases from isolation during an outbreak will have minimal effect on the final size of an outbreak, while reducing the overall number of people who are required to isolate. We use the model to quantify the dynamic viral load distribution of agents throughout an unmitigated epidemic before using it to evaluate and compare several isolation strategies in which the decision to isolate is based upon viral load. We distill the problem down to its simplest form by considering the transmission of an SIR-like pathogen in which each infected individual’s viral load is modelled using the target cell–infectious cell-virus (TIV) equations. An individual’s contagiousness is specified to be a sigmoidal function of their viral load. We account for test sensitivity as a function of viral load, and consider scenarios in which only one, or multiple tests, are available for individuals during their symptomatic illness. We evaluate isolation management strategies in which all individuals who test positive to infection are isolated, and alternative strategies in which those with a low viral load may be exempt from isolation. We compare isolation strategies in terms of the reduction in the final size of the epidemic compared to a baseline with no isolation, and how many individuals were isolated to achieve that reduction in overall infections.
@@ 2.1 Multi-scale model
We consider a fixed population of $N$ individuals, with no births or deaths occurring during the simulation period. Each individual in the population is represented as an agent $a_{k}$, where $k\in\{1,2,\cdots,N\}$. At each time step $t\in\{0,t_{\rm int},\dots,t_{\rm end}\}$, where $t_{\rm int}$ is the time step and $t_{\rm end}$ is the maximum simulation time, each agent is classified into one of three mutually exclusive sets: $\hat{S}(t)$ for susceptible, $\hat{I}(t)$ for infected and infectious, and $\hat{R}(t)$ for those who are removed and no longer infectious. The respective cardinalities of these sets are denoted by $S(t)$, $I(t)$, and $R(t)$, representing the total number of susceptible, infectious and recovered agents at time $t$.
@@
We define the contagiousness of an infectious agent (i.e., those $a_{k}\in\hat{I}$) to be the probability of infection given contact with a susceptible individual. We model contagiousness $\tau$ units of time since the time of infection, denoted by $\beta_{k}(\tau)$, as a sigmoidal function of the agent’s viral load $V_{k}(\tau)$:
@@ 51             Calculate $S(t+t_{\rm int}),\ I(t+t_{\rm int}),\ R(t+t_{\rm int})$;
While an agent’s viral load over time is a continuous trajectory, for convenience in defining and evaluating our alternative isolation strategies, we divide the infected period of each agent into four distinct phases, based on whether an agent’s viral load is low or high, and whether it is increasing or decreasing (Figure 1). The threshold viral load for classifying an agent as having a low or high viral load is denoted by $V^{*}$, and for the purposes of this paper, we choose this value such that the agent’s contagiousness $\beta=\frac{2}{3}\beta_{max}$. Combined with the sign of $dV/dt$, we compute the times of transition between the phases: $\tau_{\text{inc}}^{*}$, $\tau_{\text{peak}}$, and $\tau_{\text{dec}}^{*}$.
@@ 2.2 Testing–isolation strategies
In order for an infected individual to be isolated, they must be identified through surveillance (i.e., be recorded as a case). We only consider case detection due to symptom-based testing and do not model contact tracing, testing or quarantine of contacts, nor other aspects of surveillance such as mass screening. In our model, the probability that an infected agent is identified as a case depends on the presence of symptoms, the probability $P(\rm Tested)$ per unit time of seeking a test (symptom-dependent), and the sensitivity of the test, which is assumed to vary with viral load.
@@
Infected agents identified as cases and subject to isolation (strategy dependent) are assumed to be isolated immediately and for a fixed duration $d_{\text{iso}}$, and do not contribute to the force of infection in the population (i.e., their contagiousness is set to zero) during their isolation period. If an agent is released from isolation while still classified as infectious, they will once again contribute to the force of infection.
For each of these strategies, we consider alternative testing scenarios in which each agent may test only once, or in which multiple (unlimited) tests may be conducted. Under the single test strategy, each infected agent has only one opportunity for testing (regardless of test outcome). Therefore, an infectious agent with a low increasing viral load, if tested and positive (i.e., identified as a case), will not be isolated under the Isolate high strategy, even though they will soon become highly contagious (entering their high increasing followed by high decreasing infection phase). Under a multiple testing scenario, such an agent could test positive later in their infection and be isolated. The same outcomes for an agent can potentially play out under the Adapativen strategy at certain times in the epidemic, after the switch to only isolating cases with a high viral load has been made.
We denote by $\hat{Q}_{s}(t)$ the set of agents that are isolated at time $t$ under strategy
@@ 2.3 Strategy evaluation
Finally, to evaluate the performance of the Adaptiven strategy in more detail, we introduce an additional metric, the relative cost, $\psi_{n}$ for the Adaptiven strategy. At best, the Adaptiven strategy should perform as well as the Isolate all strategy and at worst, perform as poorly as the Isolate high strategy. We define $\psi_{n}$ as the proportionate loss of benefit (i.e., the relative cost) of the Adaptiven strategy between these two extremes:
@@
Accordingly, $\psi_{n}\to 0$ indicates that the effectiveness of the Adaptiven strategy is the same as for the Isolate all strategy (i.e., a cost of zero), while $\psi_{n}\to 1$ indicates that the Adaptiven strategy has lost (on average) all of the benefit of having isolated cases with a low viral load at the time of detection, with effectiveness degraded (on average) to the same value as for the Isolate high strategy.
@@ 2.4 Model parameterisations
All outputs presented in the main manuscript were generated with the “influenza like” model parameterisation. In the Supplementary Material, we present analagous outputs under the “SARS-CoV-2” model parameterisation.
For all scenarios presented (including those in the Supplementary Material), we consider a population of size $N=20,000$, with $\beta_{max}=0.6$, initial conditions ${S(0)=N-10}$, ${I(0)=}10$, ${R(0)=0}$, and with all other parameters specified as per the Methods and summarised in Table 1. Our choice for $\beta_{max}$ results in an epidemiological-scale basic reproduction number of $\mathcal{R}_{0}\approx 2.12$. With these parameter choices, initial conditions, and without any intervention, it is unlikely for stochastic extinction to occur during the early phase of an epidemic. The code used for this work is available at: https://github.com/Effectiveness-of-isolating-low-viral-load-Multi-scale-model.
@@ 3.1 Multi-scale dynamics
Figure 2(a) shows an exemplar realisation of the model displaying SIR-type epidemic dynamics. For the same realisation, Figure 2(b) shows the contagiousness ($\beta$) of each infectious agent on each day, differentiated by whether the agent’s contagious is increasing (purple circles, labeled $\beta^{inc}$) or decreasing (green circles, labeled $\beta^{dec}$) on that day. Prior to the peak (around day 27), many of the infected agents are in the increasing phase of their contagiousness, whereas following the peak, most agents are in the decreasing phase of their contagiousness. Probability (PDF) and cumulative (CDF) density functions for the distribution of infected agents’ contagiousness are shown in S3 Fig..
Figures 2(c)–(f) show key aspects of our multi-scale model averaged over 20 realisations. We break the infectious population down into those with low (yellow) or high (blue) viral load (Figure 2(c)), and further, as per Figure 1, into those with a low increasing, high increasing, high decreasing and low decreasing viral load (Figure 2(e)). The contribution of these subsets of the infectious population to the overall force of infection is shown in Figures 2(d) and (f). While the proportion of infected agents with a low viral load increases throughout the outbreak, their contribution to the force of infection remains small (under 20%). Furthermore, following the epidemic peak, the vast majority of those with a low viral load are in the decreasing phase of their infection (i.e., their contagiousness is low and decreasing), suggesting that an Adaptiven strategy—in which those individuals with a low viral load at the time of detection are not be required to isolate—may prove an effective and efficient policy option.
@@ 3.2 Effectiveness and efficiency of alternative testing–isolation strategies
Figure 3 shows incidence (left panels) and cumulative incidence infection (right panels) curves for a single exemplar simulation of the agent-based model under the four strategies as defined in the Methods (No isolation, Isolate high, Isolate all and Adaptive0 with the strategy switch occurring at the peak day (i.e., $n=0$)). Results are shown under a multiple test scenario (top) and single test scenario (bottom).
@@
As expected, the Isolate all strategy is the most effective strategy as all agents that test positive are isolated. Allowing for multiple tests results in nearly twice the number of cases averted compared to if only single tests are allowed (Figure 4(a) and (b)). The Adaptive0 strategy (with the switch enacted at the peak day) consistently yielded an intermediate effectiveness suggesting that, while the vast majority of those with a low viral load are in the decreasing phase of their infection following the peak, their collective contribution to transmission remains important. A more surprising result is that the efficiency—defined as the number of cases averted per person isolated—is also highest for the Isolate all strategy and lowest for the Isolate high strategy. Furthermore, efficiency is higher under the multiple tests scenario compared to the single test scenario (Figure 4((c) and (d)).
@@ 3.3 Sensitivity analyses: inclusion of asymptomatic infections and alternative within-host dynamics
To this point, we have assumed that all infected agents will display symptoms and that the probability of seeking a test per unit time increases when displaying symptoms. Our particular choice of parameters for those testing probabilities means that a substantial proportion of all infections are detected as cases, with corresponding strong impact on transmission under all strategies. We next explore a scenario in which 30% of the infected agents remain asymptomatic throughout their infectious period. Asymptomatic cases are assumed to have only a $20\%$ chance per day of being tested, fixed throughout their infectious period. This is consistent with our assumptions for pre-symptomatic infections which have a 20% chance of testing per day. Results are shown in S4 Fig.. The trends in both effectiveness and efficiency across alternative strategies remain unchanged, with the obvious effect that with 30% of the population less likely to be tested, effectiveness is lower for any given strategy.
@@ 4 Discussion
While effective at reducing transmission, NPIs, such as case isolation, can also produce unintended health and economic costs. Therefore, it is desirable to examine pandemic response policies that would enable such measures to be targeted more selectively in order to reduce their social costs while maintaining their health benefits. Here, we explored the potential for isolation decisions based on routinely collected (during a pandemic) viral load data. Specifically, we explored the hypothesis that exempting low viral load cases throughout or only during the latter stages of an outbreak could reduce the overall number of people who are isolated with minimal effect on the size of the outbreak.
@@
We made a number of simplifying assumptions in our model. A quantitative exploration of alternative isolation strategies for a particular pathogen of interest would require incorporation of additional biological and public health policy complexity. Firstly, each agent’s viral dynamics were modelled using the basic TIV model and with identical parameters (except for the initial viral load), so there is less variation across agents in our simulation than would be expected in reality. Addressing this would be technically straightforward as, at least for pathogens such as influenza and SARS-CoV-2, there are now more realistic within-host models and suitable data from which to compute population-level estimates for the joint probability distribution of within-host parameters, allowing for individual agent’s viral kinetics to be sampled. While a significant undertaking, this data-driven work would enable a more realistic population to be simulated. Secondly, our testing and isolation model is very simple. We adopted simple functional forms for both the probability of seeking a test and the probability of returning a positive result given the individual’s viral load. Empirical data on how both quantities relate to the time of exposure and symptom onset are well documented for some pathogens and could be incorporated into our multi-scale model if it were to be applied to a particular pathogen and response context of interest. Finally, delays to isolation following testing—which depend upon aspects of the public health laboratory reporting and surveillance system—are also important, but not modelled here. Again, empirical data on those delays could be incorporated into our flexible multi-scale framework.
A natural extension to our work would be to allow for a different (shorter) duration of isolation for those with low viral loads under the Adaptiven strategy. It may be that a shorter period of isolation (compared to that implemented for all cases identified prior to the peak of the epidemic) would prevent the vast majority of infection events, and result in an increase in efficiency, appropriately re-defined as the number of infections averted per person-days of isolation (rather than per person isolated).
By developing a multi-scale model of transmission dynamics, we have been able to explore the potential impacts of alternative isolation strategies on epidemic dynamics. Our results indicate that even those cases with low and decreasing viral loads make a sufficient contribution to overall transmission. Any consideration of adaptive isolation policies based on viral load would require detailed evaluation using a carefully calibrated multiscale model, for which our model could provide a foundation, and likely additional surveillance streams to ensure a net positive impact of such policies.
5 Supplementary Material